### v1

@@ Abstract
We present SpiralMLP, a novel architecture that introduces a Spiral FC layer as a replacement for the conventional Token Mixing approach. Differing from several existing MLP-based models that primarily emphasize axes, our Spiral FC layer is designed as a deformable convolution layer with spiral-like offsets. We further adapt Spiral FC into two variants: Self-Spiral FC and Cross-Spiral FC, which enable both local and global feature integration seamlessly, eliminating the need for additional processing steps. To thoroughly investigate the effectiveness of the spiral-like offsets and validate our design, we conduct ablation studies and explore optimal configurations. In empirical tests, SpiralMLP reaches state-of-the-art performance, similar to Transformers, CNNs, and other MLPs, benchmarking on ImageNet-1k, COCO and ADE20K. SpiralMLP still maintains linear computational complexity $O(HW)$ and is compatible with varying input image resolutions. Our study reveals that targeting the full receptive field is not essential for achieving high performance, instead, adopting a refined approach offers better results.
Keywords:
MLP Lightweight Vision Model
@@ 1.1 Background
Previous image classification systems primarily utilize CNN-based architectures [20, 84, 52, 56]. These systems perform exceptionally well on controlled datasets but may falter when it comes to uncontrolled bias-based situations. In recent years, Transformer-based architectures [66, 14, 27, 2] have emerged as viable alternatives to CNN-based architectures. This is largely due to their use of a self-attention mechanism that has been shown to generalize well for large datasets [53], and it also can be leveraged for downstream tasks [42]. However larger memory capacities and more regularized datasets may be required to optimize the Transformers.
In recent times, the basic MLP-based architectures [59, 39] have demonstrated promising results comparable to that of Transformers on various computer vision tasks. The dataset-efficient and lightweight architecture consists of two key components: Channel Mixing and Token Mixing. Channel Mixing entails performing feature extraction for each token by projecting feature maps along the channel dimension, while Token Mixing captures spatial information by projecting the feature map along the spatial dimension. Together, the two mixing layers enable local and global context aggregation, enhancing the robustness and decreasing the training resource consumption.
@@
In recent times, MLP-based architectures have emerged as promising alternatives to CNNs and Transformers in the realm of computer vision. The core idea behind these architectures is to leverage MLP as a fundamental building block, leading to highly efficient and effective models. The pioneering MLP-Mixer [59] processes a simple yet powerful architecture with both Token Mixing and Channel Mixing. Given a feature map $X\in\mathbb{R}^{H\times W\times C_{\text{in}}}$, where $H$, $W$ are the height and weight, $C_{\text{in}}$ is the input channel dimension, let $W^{\text{Tmix}}\in\mathbb{R}^{H\cdot W\times H\cdot W}$ denote the token mixing weight matrix, the operation applied to the preprocessed input $X^{T}\in\mathbb{R}^{C_{\text{in}}\times H\cdot W}$ is described as follows:
@@
Though MLP-Mixer provides impressive performance, it still suffers from (Eq. 1) quadratic computational complexity $O(H^{2}W^{2})$. Furthermore, using a fully-connected layer to perform token mixing makes it only accept images with fixed input size, thereby hindering its performance in the following tasks. Subsequently, the gMLP [39] proposes the spatial gating unit to enhance information integration, FNet [32] indicates token mixing can be replaced by Fourier transform, HireMLP [18] splits input into regions among which swaps elements to mimic the self-attention mechanism, ResMLP [60] employs three trainable matrices to replace LayerNorm [1], s2MLP [76] applies shifting operations, WaveMLP [58] represents pixels with complex numbers and argues that the token map is also trainable, ViP [23] applies a permutator to grasp spatial information along horizontal and vertical axes, while MorphMLP [79] modifies a gradually expands receptive field along the same axes, etc.
Nonetheless, the aforementioned MLPs have not significantly reduced the computational complexity thus leave room for SparseMLP [57], ASMLP [36], CycleMLP [6] and ATM [69]. Both SparseMLP and ASMLP perform token mixing densely along the channel dimension. In contrast, CycleMLP proposes Cycle FC with a modified yet fixed offsets $S_{H}$ and $S_{W}$ to perform a channel-wise sparser token mixing. Notably, ATM [69] designs similar yet trainable offsets to enable token mixing with a dynamic receptive field. However, these four methods perform token mixing confined within a pair of cross lines, suggesting a limited integration of feature information only along the horizontal and vertical axes.
To tackle these issues, we introduce SpiralMLP, featuring a modified fully-connected layer known as Spiral FC, which is built upon Channel FC as depicted in Fig. 1(a). Spiral FC not only ensures comprehensive awareness within the receptive field but also maintains linear computational complexity. The layout of our paper is structured as follows:
  • We provide a comprehensive description of the SpiralMLP architecture.
  • We perform experiments to highlight the superiority of SpiralMLP in comparison to other SOTAs, and conduct ablation studies to explore the functionalities of Spiral FC and discuss the optimal configuration.
  • We provide a summary of our method, coupled with insights into the prospective research to tackle the existing limitations.
@@ 2.1 Spiral FC
We aim to design a token mixing layer with the smallest possible size that still captures comprehensive spatial information within its receptive field. Upon comparing aforementioned architectures that utilize criss-cross fully-connected layers, we contend that these designs inadequately leverage the offset function. As a result, they fail to encapsulate sufficient spatial information within their receptive fields, despite having a higher parameter count.
Thus to reduce computational complexity and improve adaptability to diverse input sizes, we draw inspiration from the natural spiral shape and AttentionViz [75], where a distinctive spiral pattern is observed in transformer attention layer visualizations. Consequently, we propose the Spiral Fully-Connected Layer (abbreviated as Spiral FC) [^1] The Appendix offers a comprehensive demonstration. . This innovation aims to substitute the vanilla Token Mixing (Eq. 1) in the MLP-Mixer architecture. As illustrated in Fig. 1(b), instead of conducting convolution directly in the manner of Channel FC, the Spiral FC employs a spiral trajectory within the feature map $X\in\mathbb{R}^{H\times W\times C_{\text{in}}}$ to incorporate various features into calculations, and it is formulated as:
@@
where, $T$ is the constant period, $A(\cdot)$ is the amplitude that controls the width of the spiral trajectory, for conciseness, we formulate the amplitude function $A(\cdot)$ with the basic pattern 222 Sec. 4.1 provides additional cases with universal offset functions.:
@@
where, $A_{\text{max}}$ is the maximum amplitude. When $A_{\text{max}}=0$, the Spiral FC is identical to Channel FC, denoted as Self-Spiral FC. Conversely, when $A_{\text{max}}\neq 0$, it is termed as Cross-Spiral FC. Additionally, we employ a sliding window with a stepsize of $1$. It not only makes the Spiral FC agonistic to the input size, but also enables the flexible feature extraction through meticulously modifying the offset functions (Eqs. 4 and 5), thereby ensuring the Spiral FC operate with linear computational complexity.
@@ 3 Experiments
ConvMLP-L [34] 80.2 43 - RepMLP-Res50-g8/8 [13] 76.4 59 12.7
ConvMLP-M [34] 79.0 17 - MLPMixer-B/16 [59] 76.4 59 12.7
@@ 3.1.1 Settings
Our implementation primarily draws from DeiT [61]. The training is on a cluster of 4 NVIDIA A100 GPUs for a total of 300 epochs. The overall batch size is 512 and we employ the Top-1 accuracy for image classification.
@@ 3.1.2 Comparison with MLPs
Swin-T [42] 38.5 245.0 41.5 62.1 44.2 - - -
VanillaNet-13 [4] 74.6 397.0 41.8 62.8 44.3 - - -
CycleMLP-B5 [6] 85.9 - 42.7 63.3 45.3 24.1 46.3 57.4
HireMLP-Large (1x) [18] 105.8 424.5 44.9 - - 28.9 48.9 57.5
@@ 4 Ablation
Figure 6: (a) visualizes the effects of varying $k$ on the shape of spiral trajectory as described by Eqs. 14 and 15, while maintaining a constant $A_{\text{max}}=7$. (b) examines the effect of varying $k$ while maintaining $A_{\text{max}}$ at $3$, whereas (c) explores the influence of different $A_{\text{max}}$ values with $k$ held constant at $2$.
@@ 4.1 Update the Offset Functions
We also provide the visualizations of Eqs. 14 and 15, as depicted in Fig. 6 (a), showcasing variations with different numbers of partitions $k$.
@@ 4.2 Ablation Study on $k$
With the updated offset functions $\phi^{*}_{i}(\cdot)$ and $\phi_{j}^{*}(\cdot)$ (Eqs. 14 and 15), we evaluate the performance of variants when the number of partitions $k$ takes different values, in terms of Top-1 Accuracy on ImageNet-1k. These results are illustrated in Fig. 6 (b), with the maximum amplitude $A_{\text{max}}$ fixed at $3$. The trend reveals that the accuracy curve, in relation to $k$, initially increases, peaking at $k=2$, before diminishing.
The potential explanation of this phenomenon posits that varying $k$ influences the emphasis on the marginal regions of the receptive field, meaning that $k$ modulates the concentration of feature points at the periphery. As illustrated in Fig. 6 (a), when $k$ is set to $2$, the feature points on the edge of the receptive field are more densely clustered than when $k=4$. A too smaller $k$ value results in feature points being densely and narrowly packed, whereas a too larger $k$ scatters the feature points more widely, potentially impairing the model’s performance.
@@ 4.3 Ablation Study on $A_{\text{max}}$
We investigate several cases when the maximum amplitude $A_{\text{max}}$ takes various values. From the results shown in Fig. 6 (c), we observe an initial improvement in the Top-1 Accuracy on ImageNet-1k. However, a decline becomes evident once the $A_{\text{max}}$ exceeds $3$.
@@ 4.4 Ablation Study on Fully-Connected Layers
To illustrate the effectiveness of Spiral FC, we perform experiments on the CIFAR-10 [29] using SpiralMLP-B1444The configuration of SpiralMLP-B1 is demonstrated in Appendix. as the base architecture. In these experiments, the Spiral FC is substituted with various alternatives, including PATM from WaveMLP [58], ATMLayer from ATM [69], CycleFC from CycleMLP [6] and a RandomFC. The Random FC is architecturally identical to Spiral FC, except that the offset function is generated randomly.
SpiralFC PATM ATMLayer CycleFC RandomFC
Acc (%) 95.6 95.3 95.2 94.7 94.5
Params (M) 14 17 15 15 14
@@ 4.5 Latency Analysis
Model Params Latency Model Params Latency
(M) (ms) (M) (ms)
SpiralMLP-B4 46 47.00 CycleMLP-B4 [6] 52 57.94
SpiralMLP-B5 68 39.22 CycleMLP-B5 76 48.38
ATM-B [69] 52 64.77 PVTv2-B4 [68] 63 43.96
ATM-L 76 54.09 PVTv2-B5 82 55.71
@@ 5 Conclusion and Future Work
In this paper, we present the versatile Spiral FC, a key element of Spiral Mixing, which is intended to replace the vanilla Token Mixing. Building on this innovation, we have developed SpiralMLP, a new backbone for computer vision tasks, compatible with both PVT-style and Swin-style architectures. Among a range of experiments encompassing image classification, object detection, instance segmentation, and semantic segmentation, SpiralMLP maintains competitiveness with state-of-the-art models, achieving with a lower number of parameters, reduced computational complexity (as measured in FLOPs), and enhanced latency speed.
To the best of our knowledge, we are the first team to employ meticulously designed offset functions in order to possibly capture the most comprehensive feature information within the receptive field. This approach marks a significant difference from previous works such as CyelcMLP [6], ASMLP [36] and ATM [69], which are primarily concentrated on optimizing either static or dynamic cross-like fully connected layers.
SpiralMLP’s impressive performance necessitates further research to optimize its hyperparameters and explore a dynamic version, promising a simplified structure and more efficient information capture within its receptive field.
Appendix 0.A Related Works
0.A.1 CNN-Based
@@ 0.A.2 Transformer-Based
Appendix 0.B Compare SpiralMLP to CNNs and MHSA
@@
0.B.1 How is SpiralMLP related to CNNs?
@@
0.B.2 How is MHSA related to CNNs?
@@
0.B.3 How is SpiralMLP related to MHSA?
@@
Appendix 0.C Model Zoo and Training Details
0.C.1 Model Zoo Configurations and Performances
@@
0.C.2 Experimental Setup for Image Classification
@@
0.C.3 Experimental Setup for Object detection and Instance Segmentation
@@
0.C.4 Experimental Setup for Semantic Segmentation
@@
Table 8: Configurations of SpiralMLP variants in Swin-style.
@@
Appendix 0.D Merge Head
0.D.1 How does the Merge Head Work?
@@
0.D.2 Complexity of Merge Head.
@@
Table 9: Tensor shape of each step in Merge Head.
@@
Appendix 0.E Why Spiral FC works?
@@
Appendix 0.F Latency Analysis
We evaluate the latency speeds on one A100 across various architectures for input resolutions of $224^{2}$, $384^{2}$, and $512^{2}$. These results are detailed in Table 10. It is evident that SpiralMLP is faster than other MLP-based models when considering the scale of model size. Additionally, the PVT [67] showcases even faster processing speeds, mainly due to its utilization of dot product operations, inherently benefiting from accelerated computation. The models used for evaluation are directly extracted from their official implementations.

### v2

@@ Abstract
We present SpiralMLP, a novel architecture introduces a Spiral FC layer as a replacement for the conventional Token Mixing approach. Differing from several existing MLP-based models that primarily emphasize axes, our Spiral FC layer is designed as a deformable convolution layer with spiral-like offsets. We further adapt Spiral FC into two variants: Self-Spiral FC and Cross-Spiral FC, enabling both local and global feature integration seamlessly, eliminating the need for additional processing steps. To thoroughly investigate the effectiveness of the spiral-like offsets and validate our design, we conduct ablation studies and explore optimal configurations. In empirical tests, SpiralMLP reaches state-of-the-art performance, similar to Transformers, CNNs, and other MLPs, benchmarking on ImageNet-1k, COCO and ADE20K. SpiralMLP still maintains linear computational complexity $O(HW)$ and is compatible with varying input image resolutions. Our study reveals that targeting the full receptive field is not essential for achieving high performance, instead, adopting a refined approach offers better results.
@@ 1.1 Background
Earlier image classification systems mainly relied on CNN-based architectures [20, 84, 52, 56], which excel with controlled datasets but struggle with biased or uncontrolled conditions. Subsequently, Transformer-based architectures [66, 14, 27, 2] have emerged as alternatives, benefiting from self-attention mechanism that excel with large datasets [53] and are adaptable for various tasks [42]. However, they are often more expensive in pretraining and need specific datasets for better performance on downstream tasks.
MLP-based architectures [59, 39] have also shown promise in computer vision tasks, matching Transformer performance with a more data-efficient and lighter design. These systems use two main components: Channel Mixing, which projects features along the channel dimension, and Token Mixing, which captures spatial information by projecting feature along the spatial dimension. These mixing layers collectively enhance context aggregation, improving robustness and reducing training resource needs.
@@
The pioneering MLP-Mixer [59] proposes a simple yet powerful architecture with both Token Mixing and Channel Mixing. Given a feature map $X\in\mathbb{R}^{H\times W\times C_{\text{in}}}$, where $H$, $W$ are the height and weight, $C_{\text{in}}$ is the input channel dimension, let $W^{\text{Tmix}}\in\mathbb{R}^{H\cdot W\times H\cdot W}$ denote the token mixing weight matrix, the operation applied to the reshaped input $X^{T}\in\mathbb{R}^{C_{\text{in}}\times H\cdot W}$ is described as follows:
@@
While MLP-Mixer shows strong performance, it is limited by its quadratic computational complexity $O(H^{2}W^{2})$ (Eq. 1) and requires fixed image sizes due to its fully-connected token mixing layer. Alternatives like gMLP [39] introduces a spatial gating unit for better integration, FNet [32] uses Fourier transforms for token mixing, and HireMLP [18] mimics self-attention by swapping elements across regions. Other developments include ResMLP [60], which replaces LayerNorm with trainable matrices, s2MLP [76] utilizes shifting operations, WaveMLP [58] treats pixels as complex numbers, ViP [23] employs a permutator for spatial data, and MorphMLP [79] gradually expands its receptive field.
Despite advancements, the MLPs mentioned earlier haven’t significantly cut computational complexity, leaving an opening for SparseMLP [57], ASMLP [36], CycleMLP [6], and ATM [69]. SparseMLP and ASMLP use dense token mixing along the channel dimension, while CycleMLP introduces Cycle FC for sparser, channel-wise mixing with fixed offsets $S_{H}$ and $S_{W}$. ATM [69], on the other hand, uses trainable offsets for dynamic token mixing. However, these models restrict token mixing to horizontal and vertical axes, limiting their ability to integrate feature information across different spatial dimensions.
To address these challenges, we present SpiralMLP with its core component, Spiral FC, based on Channel FC as shown in Fig. 1(a). Spiral FC offers sufficient receptive field coverage while maintaining linear computational complexity. The paper is organized as follows:
  • We introduce the SpiralMLP architecture and its foundational Spiral FC layer.
  • We conduct experiments to demonstrate SpiralMLP’s superiority over other state-of-the-art models.
  • We perform ablation studies to explore optimal configurations, followed by conclusions and discussions on potential future improvements.
@@ 2.1 Spiral FC
We aim to design a compact token mixing layer that captures spatial information efficiently. Our review indicates that traditional designs with criss-cross fully-connected layers fail to optimize the offset function, resulting in inadequate spatial coverage. To address these challenges, we draw inspiration from natural spiral patterns and AttentionViz [75], noted for its spiral patterns in transformer attention visualizations.
As a result, we introduce the Spiral Fully-Connected Layer (Spiral FC), intended to replace standard Token Mixing (Eq. 1) in the MLP-Mixer architecture. Described in Fig. 1(b), Spiral FC leverages a spiral trajectory across the feature map $X\in\mathbb{R}^{H\times W\times C_{\text{in}}}$:
@@
where, $T$ is the constant period, $A(\cdot)$ is the amplitude that controls the width of the spiral trajectory, for conciseness, we formulate the amplitude function $A(\cdot)$ with the basic pattern 111 Sec. 4.1 provides additional cases with universal offset functions.:
@@
where, $A_{\text{max}}$ is the maximum amplitude. When $A_{\text{max}}=0$, the Spiral FC is identical to Channel FC, denoted as Self-Spiral FC. Conversely, when $A_{\text{max}}\neq 0$, it is termed as Cross-Spiral FC. Additionally, we employ a sliding window with a stepsize of $1$. It not only makes the Spiral FC agnostic to the input size, but also enables the flexible feature extraction through meticulously modifying the offset functions (Eqs. 4 and 5), thereby ensuring the Spiral FC operate with linear computational complexity.
@@ 3 Experiments
ConvMLP-L [34] 80.2 43 9.9 RepMLP-Res50-g8/8 [13] 76.4 59 12.7
ConvMLP-M [34] 79.0 17 4.0 MLPMixer-B/16 [59] 76.4 59 12.7
@@ 3.1.1 Settings
Our implementation primarily draws from DeiT [61]. The training is 4 NVIDIA A100 GPUs for a total of 300 epochs. The overall batch size is 512 and we employ the Top-1 accuracy for image classification.
@@ 3.1.2 Comparison with MLPs
CycleMLP-B5 [6] 85.9 360.3 42.7 63.3 45.3 24.1 46.3 57.4
@@ 4 Ablation
Figure 6: Visualization of varying $k$ on spiral trajectory as described by Eqs. 14 and 15, while maintaining a constant $A_{\text{max}}=3$.
Case 1 $k$
$A_{\text{max}}=3$ 1 2 3 4 5
Acc(%) 83.9 84.0 83.8 83.6 83.3
Case 2 $A_{\text{max}}$
$k=2$ 2 3 4 5 6
Acc(%) 83.8 84.0 83.7 83.4 82.9
Table 5: Experiments on $k$ and $A_{\text{max}}$. After reaching their respective peaks, both trends show a rapid decline.
@@ 4.1 Update the Offset Functions
We also provide the visualizations of Eqs. 14 and 15, as depicted in Fig. 6, showcasing variations with different numbers of partitions $k$.
SpiralFC PATM ATMLayer CycleFC RandomFC
Acc (%) 95.6 95.3 95.2 94.7 94.5
Params (M) 14 17 15 15 14
@@ 4.2 Ablation Study on $k$
We updated the offset functions $\phi^{*}_{i}(\cdot)$ and $\phi_{j}^{*}(\cdot)$ (Eqs. 14 and 15) to analyze how varying the number of partitions $k$ affects Top-1 Accuracy on ImageNet-1k. Results, shown in Tab. 5 with a constant maximum amplitude $A_{\text{max}}$ of $3$, indicate that accuracy initially rises, peaks at $k=2$, then decreases.
This trend suggests that different $k$ values alter the focus on the peripheral regions of the receptive field, where $k=2$ results in denser clustering of feature points along the edges compared to $k=4$, as seen in Fig. 6. Lower $k$ values cause a dense, narrow concentration of features, while higher values disperse them too widely, potentially reducing model effectiveness.
@@ 4.3 Ablation Study on $A_{\text{max}}$
We investigate several cases when the maximum amplitude $A_{\text{max}}$ takes various values. From the results shown in Tab. 5, we observe an initial improvement in the Top-1 Accuracy on ImageNet-1k. However, a decline becomes evident once the $A_{\text{max}}$ exceeds $3$.
@@ 4.4 Ablation Study on Fully-Connected Layers
To illustrate the effectiveness of Spiral FC, we perform experiments on the CIFAR-10 [29] using SpiralMLP-B1333The configuration of SpiralMLP-B1 is demonstrated in Appendix. as the base architecture. In these experiments, the Spiral FC is substituted with various alternatives, including PATM from WaveMLP [58], ATMLayer from ATM [69], CycleFC from CycleMLP [6] and a RandomFC. The Random FC is architecturally identical to Spiral FC, except that the offset function is generated randomly.
@@ 4.5 Latency Analysis
Model Params(M) Latency(ms)
SpiralMLP-B4 46 47.00
SpiralMLP-B5 68 39.22
CycleMLP-B4 [6] 52 57.94
CycleMLP-B5 76 48.38
ATM-B [69] 52 64.77
ATM-L 76 54.09
PVTv2-B4 [68] 63 43.96
PVTv2-B5 82 55.71
@@ 5 Conclusion and Future Work
In this paper, we present Spiral FC, part of Spiral Mixing designed to replace traditional Token Mixing. We introduce SpiralMLP, a new computer vision framework compatible with PVT-style and Swin-style architectures. SpiralMLP performs comparably to leading models while using fewer parameters and less computational power.
We believe we are the first to use carefully designed offset functions to capture comprehensive feature information, setting us apart from models like CyelcMLP [6], ASMLP [36] and ATM [69], which focus on optimizing cross-like layers. Given its strong performance, further research into optimizing SpiralMLP’s hyperparameters could lead to even more efficient information capture.
Appendix A Related Works
A.1 CNN-Based
@@ A.2 Transformer-Based
Appendix B Compare SpiralMLP to CNNs and MHSA
@@
B.1 How is SpiralMLP related to CNNs?
@@
B.2 How is MHSA related to CNNs?
@@
B.3 How is SpiralMLP related to MHSA?
@@
Table 8: Configurations of SpiralMLP variants in PVT-style.
Appendix C Model Zoo and Training Details
C.1 Model Zoo Configurations and Performances
@@
C.2 Experimental Setup for Image Classification
@@
C.3 Experimental Setup for Object detection and Instance Segmentation
@@
C.4 Experimental Setup for Semantic Segmentation
@@
Table 9: Configurations of SpiralMLP variants in Swin-style.
@@
Appendix D Merge Head
D.1 How does the Merge Head Work?
@@
D.2 Complexity of Merge Head.
@@
Table 10: Tensor shape of each step in Merge Head.
Appendix E Why Spiral FC works?
@@
Model Params(M) Throughput (/sec) Wall-Clock (s)
Spiral-B1 14 425 22.8
Spiral-B2 24 253 30.5
Spiral-B3 34 143 52.2
Spiral-B4 46 94 72.5
Spiral-B5 68 89 79.2
Cycle-B1 [6] 15 347 23.8
Cycle-B2 27 207 36.0
Cycle-B3 38 117 60.4
Cycle-B4 52 79 84.0
Cycle-B5 76 75 92.0
ResNet-152 [20] 60 152 45.7
Vit-B/16 [14] 86 102 75.2
Deit-B/16 [61] 86 97 78.3
Table 11: Throughput and wall-clock test.
Appendix F Latency & Throughput Analysis
As shown in Tab. 11, throughput testing is conducted on a single NVIDIA A100 GPU, with a batch size of 32 and image resolution of 3x224x224; wall-clock testing is conducted on ImageNet-1k with a single NVIDIA A100 GPU, with a batch size of 64, the number of total batches to 100 and image resolution of 3x224x224. Furthermore, we evaluate the latency speeds on one A100 across various architectures for input resolutions of $224^{2}$, $384^{2}$, and $512^{2}$. These results are detailed in  Tab. 12.
It is evident that SpiralMLP is faster than other MLP-based models when considering the scale of model size. Additionally, the PVT [67] showcases even faster processing speeds, mainly due to its utilization of dot product operations, inherently benefiting from accelerated computation. The models used for evaluation are directly extracted from their official implementations.